OpenStudy (anonymous):

What is the equation of the line that passes through (4,2) and parallel to 3x-2y=-6?

OpenStudy (anonymous):

that's wrong cuz that's not one of the choices

hero (hero):

I overlooked something, sorry

hero (hero):

Oky, I'm messing this up pretty badly

hero (hero):

One last try

OpenStudy (anonymous):

ok

hero (hero):

3x-2y=-6 y = 3x/2 + 3 original equation in slope intercept form y - 2 = 3/2(x - 4) y = 3x/2 - 6 + 2 y = 3x/2 - 4

hero (hero):

That has to be right

OpenStudy (anonymous):

thank u